# Identifying Protein Co-regulatory Network Logic by Solving B-SAT Problems through Gate-based Quantum Computing

### v1

@@ I Introduction
In this paper, we explore the use of quantum computing applied to a central challenge in computational biology, specifically reverse-engineering of regulatory logic networks. We use, as our model, a well-studied biological regulatory network of morphogens/signaling proteins that are expressed during brain development and that help establish the expression patterns of transcription factors (TFs). Specifically, Fgf8 is a signaling protein essential in establishing the midbrain-hindbrain boundary [9][10]. It controls the graded expression of TFs such as COUP-TFI, Emx2, Pax6, and Sp8 that help regulate expressions of genes whose gene products (i.e. proteins) convey positional or area identities to cortical progenitors [11]. We use the corresponding regulatory logic model proposed by [12] to generate simulated time course instances then attempt to use quantum algorithms to reverse engineer the original regulatory logic from these simulated data points . In keeping with our previous work, this problem was articulated as a Boolean Satisfiability problem (B-SAT) to facilitate the design of a corresponding quantum circuit. Specifically, the regulatory logic governing state transition of each network node was stated as a Boolean expression where the solution consists of a set of logical operators and upstream regulators that render the expression True. At its core, SAT remains a search problem, where the choice of an expression $f(s_{1},s_{2},...,s_{n})$ must match specific instances. Grover’s algorithm is framed in this exact way, with an arbitrary function enclosed in an ’oracle’ and a search space of $2n$ possible values. Using Grover’s algorithm and the principles of quantum superposition of states, the complexity of the search was reduced to $\mathcal{O}(\sqrt{n})$ and the expression corresponds to a logic statement governing state transition. To solve this logic estimation problem, we leveraged the implementation of Grover’s algorithm and the oracle synthesizer available under the Qiskit Python programming environment [13] to generate the corresponding quantum circuit. Further reductions in the complexity of the quantum circuit produced by the built-in Qiskit oracle synthesizer were obtained by carefully introducing additional problem-specific information. This optimized circuit was used to recover known regulatory logic with the populations of solutions demonstrating the contribution of intrinsic noise and its interdependency with the number of reference constraints imposed by the simulated experimental observations.
@@ II-B Assumptions and General Formulation of State Equations
Parametrizing the problem in this way effectively limits the complexity of modelling a $K$-variable network to an upper bound of $2K$ as opposed to the potentially much larger number of parameters necessary to fully define the general conjunctive normal form in which literals can appear more than once and in multiple different clauses. Though each clause has been reduced to a single literal, their combined use is sufficient to generate all of the state transition equations presented in Weidner et al. [14]. In the case of more complex state transition dynamics where certain literals are suspected to be clauses of several literals instead of one, the latter can be formulated as component BSAT problems and solved independently. Importantly, deployment of this problem on current Noisy Intermediate Scale Quantum (NISQ) platforms implicitly requires that certain simplifying assumptions be made to accommodate the available quantum volume of the device, given in terms of the maximum quantum circuit width (number of qubits) and depth (number of gates). A more complex model form describing network state transition could be applied to this problem; that said, this compact representation not only adequately captures the dynamics of the system but also facilitates deployment on current NISQ hardware given current limitations.
@@ II-E Implementing the Quantum Circuit with the Qiskit SDK
The second implementation leverages a built-in high-level library function in IBM’s Qiskit SDK which converts a high-level symbolic equation like the one in Section II-B into a Grover oracle circuit directly. Leveraging the high-level functionality is quite straightforward: we need only define the general state transition logic equation for our specific model and given sample data, then use the PhaseOracle() function from qiskit.circuit.library and apply it to our Boolean expression containing the provided sample constraints. Optionally, a separate Boolean optimization library such as sympy can be used to simplify the expression prior to passing it to the phase oracle generator. Implementation using this high-level function has the advantage of using fewer qubits since it omits the use of ancillaries, sticking more closely to the “all-in-place” style shown in most textbook representations of Grover’s algorithm. In brief, manual assembly is much more flexible and open to customization, leaving more circuit volume in which to optimize the configuration. However, this approach is also more time consuming and requires more attention to detail to be implemented correctly. In comparison, the function-assisted implementation currently executes with a faster runtime but offers virtually very little opportunity for additional optimization.
@@ III Results
In the case of this example problem, both the manual circuit designed using low-level functions and the design assisted by high-level functions (Figure 1) produce the same results under noise-free quantum simulator conditions (Figure 2), consistent with expectations. Accordingly, we will discuss the output produced using the high-level functions and compare runtime performance between both methods.
@@ III-A Execution in a Noiseless Quantum Simulation Environment
First, if we consider 8 network state profiles as constraints and attempt to recover the regulatory logic governing the expression of protein Fgf8 in the noiseless quantum simulator environment, we obtain the frequency distribution of solutions shown in Figure 3 and Table I. We then filter these solutions to select for those proposed state transition models that comply with all 8 data-imposed constraints and convert each of these bitstring outputs into the Boolean expression representation. This conversion produces a total of 51 possible logical state transition models for Fgf8 satisfying the given constraints, all of which are valid for the given (sparse) dataset. Though not the top-ranking solution, we do recover the exact original logic proposed by Weidner et al. 2023 [14] among the most probable (2.04% occurrence) (model 14). Nonetheless, many other possible statements are also capable of satisfying our sparse constraints, with these occurring with individual probabilities of  2%.
@@ III-B Execution on Real-World NISQ Hardware
In the previous section, our multiple experiments verified that our algorithm can recover the true state transition logic for protein Fgf8 in an ideal noiseless simulated quantum environment. However, the key advantage of quantum computing resides in the potential for an exponential speedup offered by a sufficiently large system of noise-corrected physical qubits, at a scale exceeding classical simulation of such systems. We conducted numerical experiments using both low-level and high-level function implementations of our circuits deployed on IBM’s Eagle r3 quantum chip with 10,000 shots to characterize how real-world noise might affect our results. Starting with our benchmark problem using 8 network state profiles as constraints we obtain the outputs shown in Figure 5 and Figure 6 respectively.
@@ IV Discussion
In this work, we present promising results showing that it is possible to harness the scalability of quantum computing to reverse engineer the regulatory logic of a biological network. This a very central problem as we begin to understanding the importance of network topology and dynamics in contributing to persistent illness and treatment resistance. Using the example of a well-documented regulatory circuit describing the control of local developmental cues in the mammalian cortex, we recover the logic rules governing its behavior by formulating a Boolean satisfiability problem that can be deployed as a quantum circuit and solved using Grover’s algorithm. Applying high-level functions that optimally leverages the basic features of quantum circuitry delivered more better discriminated and higher quality solutions but was less efficient. While runtime is heavily dependent on the QPU technology, recovering logic statements satisfying 8 network state constraints on the IBM Eagle r3 system required a typical runtime of $\sim$3m:00s using the high-level function design, compared with $\sim$1:00s using low-level functions. Moreover, our implementation based on low-level functions resulted in a lower instruction count on QPU (100568 vs 226911), likely requiring higher qubit usage and associated increases in noise. While more susceptible to noise, our use of low-level functions would also be more amenable to broader platform-specific tuning. In contrast, the oracle generated using high-level functions would be much less flexible.
Our implementation exactly recovered the true logic statement for the specific protein of interest as well as a number of competing models that supported the few network states used as constraints. An exhaustive list of feasible solutions was generated by conventional means and these exactly matched our quantum results obtained with the noise-free quantum simulator. In this regard, it is important to note that the appearance of multiple competing solutions when solving for 8 expression profile constraints should not be interpreted as noise but instead is a direct consequence of this problem being under-constrained mathematically. Nonetheless, implementation of this circuit on a real-world quantum platform highlighted some of the limitations of current NISQ hardware. For example, gate depths in the order of a few thousand approached hardware limitations beyond which the output would consist primarily of noise. This paper did not focus on optimizing the algorithm for a specific QPU’s topology, nor optimizing the core circuit operations significantly. Future work would involve such optimizations, and this would almost certainly expand the utility of this algorithm running on the same hardware.
Importantly, the potential of scale was not explored in this example. The parameterization used to model this small 5-protein network was relatively simple, consisting of a product of candidate state variables with an optional inversion. Systems with more complex network dynamics might be addressed using additional parameters and would in turn offer a better test of this technology. For example, instead of the simple logic $A\land B\land C\land D…$ one could allow for $(A\lor B)\land C\ \land D$ or similar. One could also allow for a broader variety of operations, such as NAND, XOR, etc… to describe more complex state transition logic. Such increases in complexity may better illustrate the benefit over classical algorithms and better delineate the threshold of “quantum advantage” in problems such as these.

### v2

@@ I Introduction
In this work, our primary focus is to explore the practical feasibility of applying quantum computing in the field to reverse engineer immune response networks at a scale that is immediately useful in advancing the network-informed design of immune modulatory interventions, both preventative and restorative. Towards this, we use as our model a well-studied biological regulatory network of morphogens/signaling proteins that are expressed during brain development and that help establish the expression patterns of transcription factors (TFs). Specifically, Fgf8 is a signaling protein essential in establishing the midbrain-hindbrain boundary [9][10]. It controls the graded expression of TFs such as COUP-TFI, Emx2, Pax6, and Sp8 that help regulate expressions of genes whose gene products (i.e. proteins) convey positional or area identities to cortical progenitors [11]. We use the corresponding regulatory logic model proposed by [12] to generate simulated time course instances then attempt to use quantum algorithms to reverse engineer the original regulatory logic from these simulated data points . In keeping with our previous work, this problem was articulated as a Boolean Satisfiability problem (B-SAT) to facilitate the design of a corresponding quantum circuit. Specifically, the regulatory logic governing state transition of each network node was stated as a Boolean expression where the solution consists of a set of logical operators and upstream regulators that render the expression True. At its core, SAT remains a search problem, where the choice of an expression $f(s_{1},s_{2},...,s_{n})$ must match specific instances. Grover’s algorithm is framed in this exact way, with an arbitrary function enclosed in an ’oracle’ and a search space of $2n$ possible values. Using Grover’s algorithm and the principles of quantum superposition of states, the complexity of the search was reduced to $\mathcal{O}(\sqrt{n})$ and the expression corresponds to a logic statement governing state transition. To solve this logic estimation problem, we leveraged the implementation of Grover’s algorithm and the oracle synthesizer available under the Qiskit Python programming environment [13] to generate the corresponding quantum circuit. Further reductions in the complexity of the quantum circuit produced by the built-in Qiskit oracle synthesizer were obtained by carefully introducing additional problem-specific information. We show that both the manually designed circuit assembled using low-level functions and the optimized circuit leveraging high-level functions correctly recover the true logic in a noiseless quantum simulation environment when determinacy is supported. As biomedical data is invariably time-consuming, resource intensive, invasive and often risky to collect, the reverse engineering of these biological networks from very sparse data is of primary importance. Interestingly, when presented with an underdetermined problem where multiple competing solutions legitimately exist, we find that actual NISQ hardware recovers the true logic with a higher ranking than that obtained under noiseless conditions. Likewise, we also find that the use of low-level functions resulting in a lower instruction count but increased qubit usage and potentially higher noise may deliver a more biologically relevant ranking highlighting the potentially useful role for noise inherent in these platforms to provide improved sampling of plausible solutions.
@@ II-B Assumptions and General Formulation of State Equations
Parametrizing the problem in this way effectively limits the complexity of modelling a $K$-variable network to an upper bound of $2K$ as opposed to the potentially much larger number of parameters necessary to fully define the general conjunctive normal form in which literals can appear more than once and in multiple different clauses. Though each clause has been reduced to a single literal, their combined use is sufficient to generate all of the state transition equations presented in [14]. In the case of more complex state transition dynamics where certain literals are suspected to be clauses of several literals instead of one, the latter can be formulated as component BSAT problems and solved independently. Importantly, deployment of this problem on current Noisy Intermediate Scale Quantum (NISQ) platforms implicitly requires that certain simplifying assumptions be made to accommodate the available quantum volume of the device, given in terms of the maximum quantum circuit width (number of qubits) and depth (number of gates). A more complex model form describing network state transition could be applied to this problem; that said, this compact representation not only adequately captures the dynamics of the system but also facilitates deployment on current NISQ hardware given current limitations.
@@ II-E Implementing the Quantum Circuit with the Qiskit SDK
The second implementation leverages a built-in high-level library function in IBM’s Qiskit SDK which converts a high-level symbolic equation like the one in Section II-B into a Grover oracle circuit directly. Leveraging the high-level functionality is quite straightforward: we need only define the general state transition logic equation for our specific model and given sample data, then use the PhaseOracle() function from qiskit.circuit.library and apply it to our Boolean expression containing the provided sample constraints. Optionally, a separate Boolean optimization library such as sympy can be used to simplify the expression prior to passing it to the phase oracle generator. Implementation using this high-level function has the advantage of using fewer qubits since it omits the use of ancillaries, adhering more closely to the “all-in-place” style shown in most textbook representations of Grover’s algorithm. In brief, manual assembly is much more flexible and open to customization, leaving more circuit volume in which to optimize the configuration. However, this approach is also more time consuming and requires more attention to detail to be implemented correctly. In comparison, the function-assisted implementation currently executes with a faster runtime but offers virtually very little opportunity for additional optimization.
@@ III Results
In the case of this example problem, both the manual circuit designed using low-level functions and the design assisted by high-level functions (Figure 1) produce the same results under noise-free quantum simulator conditions (Figure 2), consistent with expectations. Accordingly, we shall discuss the output produced using the high-level functions and compare runtime performance between both methods.
@@ III-A Execution in a Noiseless Quantum Simulation Environment
First, if we consider 8 network state profiles as constraints and attempt to recover the regulatory logic governing the expression of protein Fgf8 in the noiseless quantum simulator environment, we obtain the frequency distribution of solutions shown in Figure 3 and Table I. We then filter these solutions to select for those proposed state transition models that comply with all 8 data-imposed constraints and convert each of these bitstring outputs into the Boolean expression representation. This conversion produces a total of 51 possible logical state transition models for Fgf8 satisfying the given constraints, all of which are valid for the given (sparse) dataset. Though not the top-ranking solution, we do recover the exact original logic proposed by [14] among the most probable (2.04% occurrence) (model 14). Nonetheless, many other possible statements are also capable of satisfying our sparse constraints, with these occurring with individual probabilities of  2%.
@@ III-B Execution on Real-World NISQ Hardware
In the previous section, our multiple experiments verified that our algorithm can recover the true state transition logic for protein Fgf8 in an ideal noiseless simulated quantum environment. However, the key advantage of quantum computing resides in the potential for an exponential speedup offered by a sufficiently large system of noise-corrected physical qubits, at a scale exceeding classical simulation of such systems. We conducted numerical experiments using both low-level and high-level function implementations of our circuits deployed on IBM’s Eagle r3 quantum chip with 10,000 shots to characterize how real-world noise might affect our results. Starting with our benchmark problem using 8 network state profiles as constraints we obtain the outputs shown in Figure 5 and Figure 6 respectively. S
@@ IV Discussion
In this work, we present promising results showing that it is possible to harness the scalability of quantum computing to reverse engineer the regulatory logic of a biological network. This a very central problem as we begin to understanding the importance of network topology and dynamics in contributing to persistent illness and treatment resistance. Using the example of a well-documented regulatory circuit describing the control of local developmental cues in the mammalian cortex, we recover the logic rules governing its behavior by formulating a Boolean satisfiability problem that can be deployed as a quantum circuit and solved using Grover’s algorithm. As mentioned previously, the recovery of feasible solutions from very sparse data is of broad interest in biomedical research. Our implementation exactly recovered the true logic statement for the specific protein of interest as well as a number of competing models that supported the few network states used as constraints. An exhaustive list of feasible solutions was generated by conventional means and these exactly matched our quantum results obtained with the noise-free quantum simulator. While the formulation of this binary constraint satisfaction problem as two quantum circuit variants cannot and does not fundamentally resolve the indeterminacy of the sparse data problem (i.e. 8 data points), it is interesting to observe that the migration to the actual NISQ hardware and the use of low level functions, both contributors to increased noise, appear to support a broader search of the solutions space resulting in a better ranking of the true logic statement. The contribution of noise in improving search results is not unfamiliar. Indeed, the noise inherent to an NISQ environment may confer some of the same advantages of a stochastic local search [16] but where this noise is embedded directly into the parameter space [17]. Indeed, noise as a feature to be leveraged rather than simply mitigated has been gaining attention in the quantum community [18, 19].
@@
In terms of execution performance, circuit design optimization using high-level functions, while promoting noise mitigation, resulted not only in less parsimonious solutions and poorer ranking but also performed less efficiently. While runtime is heavily dependent on the QPU technology, recovering logic statements satisfying 8 network state constraints on the IBM Eagle r3 system required a typical runtime of $\sim$3m:00s using the high-level function design, compared with $\sim$1:00s using low-level functions. Moreover, our implementation based on low-level functions resulted in a lower instruction count on QPU (100568 vs 226911), likely requiring higher qubit usage and associated increases in noise. While more susceptible to noise, our use of low-level functions would also be more amenable to broader platform-specific tuning. In contrast, the oracle generated using high-level functions would be much less flexible.
Despite the transportability of this simple binary logic problem to a much broader range of use cases, this analysis remains preliminary, and work is ongoing to investigate the robustness of this approach to the choice of initial data set as well as its scalability with increasing network complexity of the network and the corresponding increase in the number of constraints. In the current problem formulation, a larger number of samples translates into increasing gate depth. Indeed, given the limitations of current NISQ hardware gate depths in the order of a few thousand approach a threshold beyond which the output would consist primarily of noise. Mitigation of decoherence [20] and optimization of resource usage as with increasing gate depth are active areas of research with optimization strategies in quantum machine learning ML in large data [21] being especially important as we continue to evaluate this aspect. It is important to recall however that we are especially focused on the identification of network logic in data poor situations making our primary challenge that of network complexity. In this regard, the scale of the reverse engineering problem in this case is directly dependent on the in-degree or number of upstream regulators of a given target node (e.g. Fgf8) since the transition logic statement for each target may be resolved individually. In this example problem the target Emx2 presented with the highest in-degree, namely 5 upstream regulators. Though the average in-degree is typically $<8$ [20], a small fraction of highly central target genes in transcriptional regulatory networks may be mediated by as many as 1,000 upstream transcription factors [23]. While new quantum chip architectures emerge at a rapid pace, a possible approach for deployment on current hardware could consist in simply sub-dividing a highly regulated gene into smaller functional subunits. In addition to the moderate degree of connectivity, the parameterization used to model this small 5-protein network was relatively simple, consisting of a product of candidate state variables with an optional inversion. Systems with more complex network dynamics might be addressed using additional parameters and would in turn offer a better test of this technology. For example, instead of the simple logic $A\land B\land C\land D…$ one could allow for $(A\lor B)\land C\ \land D$ or similar. One could also allow for a broader variety of operations, such as NAND, XOR, etc… to describe more complex state transition logic. Such increases in complexity may better illustrate the benefit over classical algorithms and better delineate the threshold of “quantum advantage” in problems such as these.
Finally, the scope of this paper did not include additional circuit optimization to a specific QPU’s topology, nor were the core circuit operations optimized significantly. Future work would involve such hardware-specific optimizations. Moreover, while we did not conduct extensive comparisons with performance classical computational methods in this analysis it is important to recall that the primary focus of this first investigation was feasibility. Indeed, the main objective of this analysis was to explore the practicality in the biomedical field of applying quantum computing to the reverse engineering of immune response networks at a scale that is immediately useful in advancing the network-informed design of immune interventions, both preventative and restorative. In exploring this, we found that noise inherent to quantum platforms may promote a more pro-active exploration of the optimization space and deliver better answers. Even in a worst-case scenario where quantum performance were equivalent, the potential for more biologically plausible solutions is without question a highly impactful contribution.